OpenStudy (anonymous):

(x^2 + 3x + 1)(x^2 + x + 2)

OpenStudy (anonymous):

write the answer in descending powers of X

OpenStudy (anonymous):

\[x^4+4 x^3+6 x^2+7 x+2 \]

OpenStudy (anonymous):

@robtobey ... explain your answer.

OpenStudy (anonymous):

\[x^2\left(x^2+3 x+1\right) +3 x\left(x^2+3 x+1\right) +2 \left(x^2+3 x+1\right) \]\[x^4+3 x^3+x^2+3 x^3+9 x^2+3 x+2 x^2+6 x+2 \]\[x^4+6 x^3+12 x^2+9 x+2 \]Actually I type in the following instruction to Mathmatica\[\left(x^2+x+2\right) \left(x^2+3 x+1\right)\text{//}\text{Expand} \]and posted the result.

OpenStudy (anonymous):

Awesomeeee !!!!! @robtobey

OpenStudy (anonymous):

Sorry. type should have been spelled, typed.